OpenStudy (anonymous):

The Fiedler family has up to $130,000 to invest. They decide that they want to have at least $40,000 invested in stable bonds yielding 5.5% and that no more than $60,000 should be invested in more volatile bonds yielding 11%. How much should they invest in each type of bond to maximize income if the amount in the stable bond should not exceed the amount in the more volatile bond? What is the maximum income? In the volatile bonds = not > than 60,000 $

OpenStudy (anonymous):

In the volatile bonds = not > than 60,000 $ To get maximum yield , you have to invest 60,000 $ in 11% bonds So amount invested in stable bonds = 130.000 - 60,000 = 70,000 $ This is more than 60,000 $ Both the conditions given cannot be satisfied.

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

u tag i dont know much people

OpenStudy (anonymous):

@mathmath333 @Jhannybean @hartnn

OpenStudy (anonymous):

@hartnn last question

OpenStudy (anonymous):

again? @hartnn

OpenStudy (anonymous):

ok

OpenStudy (aric200):

Previously posted here. http://openstudy.com/study#/updates/51a02c32e4b04449b22201c1

OpenStudy (anonymous):

correct? plz check vertices plz

OpenStudy (aric200):

This is an Exam I assume So I am not gonna help you more than I already did.

OpenStudy (anonymous):

@aric i already work it out but my teacher said do it again and said:Check the vertices again. Use the controlling parameters on the dimensions to be tiled as 2 constraints

OpenStudy (aric200):

Show me your previous work

OpenStudy (anonymous):

In the volatile bonds = not > than 60,000 $ To get maximum yield , you have to invest 60,000 $ in 11% bonds So amount invested in stable bonds = 130.000 - 60,000 = 70,000 $ This is more than 60,000 $ Both the conditions given cannot be satisfied.

OpenStudy (anonymous):

@aric200 but my teacher said check again

OpenStudy (anonymous):

plz explain

OpenStudy (anonymous):

@hartnn plz help dear i am waiting since an hour

OpenStudy (anonymous):

@ganeshie8 @Abhisar

OpenStudy (anonymous):

@Compassionate @Hero @Ashleyisakitty @AlexisPooBear

OpenStudy (mathmath333):

Let \("x"\) be the number of dollars invested in the "stable bonds" let \("y"\) be the number of dollars invested in the "more volatile bonds" maximum income \(P=1.055x+1.11y\) u have to maximize \(P\) so graph the following inequalities \(\large\tt \begin{align} \color{black}{x\geq 40000\\~\\ y\leq 60000\\~\\ x\leq y\\~\\ x+y\leq 130000\\~\\ }\end{align}\) note their edge points and find the maximum value for \(P\)

OpenStudy (anonymous):

@mathmath333 thanks finally

OpenStudy (anonymous):

40000+60000<=130000 100000 <=130000 @mathmath333

OpenStudy (mathmath333):

u have to graph the inequalities

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i use desmos and its not coming in desmos plzz help @mathmath333

OpenStudy (anonymous):

P= 1.055(40000)+1.11(60000) P=50200+66600 P=116800

OpenStudy (anonymous):

Check plzz @mathmath333

OpenStudy (mathmath333):

https://www.desmos.com/calculator/avuzhpu5yh

OpenStudy (mathmath333):

the edge points here are \(\large\tt \begin{align} \color{black}{(40000,40000)\\~\\ (40000,60000)\\~\\ (60000,60000)\\~\\ }\end{align}\)

OpenStudy (mathmath333):

see which one maximizes \(P\)

OpenStudy (anonymous):

(40000,60000)

OpenStudy (mathmath333):

i think \((60000,60000)\) will be the maximum

OpenStudy (anonymous):

@BT_92